please help me figure out steps to following problem: 5-(-4)^2 divided by (-8) x (32 -17) the answer is 35 on calculator but cant seem to figure out how to get from the problem to the answer
Answer by chessace(471)

First compute the () contents:
5-16/(-8)*(15)
Now cancel in the "fraction" by (-8) note that "* before /" convention:
= 5+(2/15) = 5.13333.
OK, apparantly your calculator treats * and / as equals and "first come first served".
5+(2*15) = 35
Look at the original form of problem and note whatever convention you've been taught. Is the (32-17) in the numerator or denominator?
I always use extra () to make sure the order of operations when / or - are present before * and + respectively.